40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4309     Entry time: Wed Feb 16 23:54:34 2011
Author: rana 
Type: Update 
Category: CDS 
Subject: Today's CDS problems 
  1. ETMX cannot load his filter coefficients. Even if I change the file, the load button doesn't work. I tried changing the lockin filters but they won't load.
  2. The ETMX filter modules appear to have 2048 for some of the modules and 16384 for some of the others. How come the make script doesn't make them all 16384?
  3. There are a bunch of kill/start scripts in the scripts/ directory instead of in the scripts/FE/ directory. Did this get reverted after a new code exchange?
  4. I tried restarting the code using c1startscx, but that doesn't work correctly. It cannot find the burtrb and burtwb files even though they are in the normal path.
  5. Kiwamu was using a bunch of cockamamie filters I found.
  6. I can't get any minute trend data. I tried on rossa, rosalba, and op440m.
MC damp dataviewer diaggui AWG c1lsc c1ioo c1sus c1iscex c1iscey RFM The Dolphins Sim.Plant Frame builder TDS Cabling
                             
ELOG V3.1.3-